OpenStudy (austinl):

Express the given complex number in the form \(\Large{R(\cos(\theta)+i\sin(\theta))=R e^{i\theta}}\) 1) -3 2) \(\sqrt{3}-i\)

OpenStudy (kmeis002):

Whats the number?

OpenStudy (kmeis002):

-3 has a radius of 3 along the - x axis. So, think of polar coordinates. On the -x axis, thats 180 deg or pi radians. So, R = 3 and \[\theta = \pi \] Therefore -3 can be rewritten as \[ 3 e^{\pi} \]

OpenStudy (kmeis002):

Correction: \[ 3 e^{i\pi} \]

OpenStudy (dan815):